if a ball strikes the ground with velocity v, having coieffecent of restitutuion as e, what is the velocity of the rebounded ball.
how to calculate the velocity of the ball using law of conservation of momentum?

Grade:11

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Hello student,
The coefficient of restitution is given by:
\text{Coefficient of Restitution } (C_{R}) = \frac{\text{Relative Speed After Collision}}{\text{Relative Speed Before Collision}}
Hence,
Relative Speed after collision would be given by:
\text{Relative Speed After Collision}=\text{Coefficient of Restitution } (C_{R}) \times {\text{Relative Speed Before Collision}}=e\times v
For finding the velocity, you can use the conservation of momentum and the coefficient of restitution.
Regards
Sumit